# Giovanni Compiani Ilya Morozov Stephan Seiler

### v2

@@ 2.3 Interpretation of Principal Components
Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect key product features and benefits. Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly. Such texts may convey functional benefits that are hard to quantify. Similarly, photos may showcase design features like color, casing style, or the presence of a keyboard. By extracting principal components from texts and images, our approach identifies these dimensions of product differentiation and uses them to capture which products are close substitutes.
Given this interpretation, researchers should not treat the coefficients on the principal components as causal in the sense that changing product images or textual descriptions does not necessarily cause a change in substitution patterns. For example, if the principal components serve as proxies for choice-relevant product attributes, these attributes need not change if the seller modifies the product’s description or visual design. Hence, the principal components should be held fixed in counterfactual analyses. For this reason, our approach is not suited for studying optimal product design and positioning, or any other questions that require causal estimates of the effects of non-price characteristics on choices.
In some markets, texts and images might not convey information relevant to consumer choices, in which case the principal components would fail to predict substitution patterns. However, we consider this scenario unlikely, as sellers have strong incentives to highlight important product attributes in titles and descriptions, and online platforms may encourage them to do so. Similarly, when writing reviews, consumers often mention product aspects they think others will find useful. Consistent with this idea, we find that in all product categories we study, at least some image- or text-based principal components predict substitution patterns.
Similar to observed product attributes in standard mixed logit demand models, we treat principal components as fixed over time. This assumption is reasonable because principal components capture time-invariant product attributes, functional benefits, and visual design features. While sellers could theoretically modify images and textual descriptions to reposition their products, we find that in practice, product images, titles, and descriptions rarely change in our empirical context (see Appendix C for details).
2.4 Advantages of Our Approach
To understand how our approach relates to existing methods, we compare it to the standard approach commonly used in demand estimation, which relies on observed product attributes. These attributes are usually incorporated into a discrete choice demand model, such as a random coefficients logit model, where substitution patterns are captured through the interaction between product attributes and preference heterogeneity.
While common, this approach presents two practical challenges. First, researchers rarely observe all attributes that differentiate products. Instead, they often rely on third-party datasets, where data providers report a limited set of attributes selected based on unknown criteria. Others gather their own data using a combination of web scraping and human coders—–a costly and time-consuming process that typically limits researchers to collecting only easily quantifiable attributes. Since attributes are selected subjectively in both cases, they may not align with those most relevant to consumer choices. Our approach avoids collecting data on product attributes and instead extracts information about product substitution from text and image data.
The second challenge is that consumers often consider visual design and functional benefits—dimensions that are difficult to capture through observed attributes. For instance, in categories like clothing or home decor, product aesthetics can strongly influence choices. A shopper selecting a dress might care about overall style, design pattern, or neckline type—details visible in photos but difficult to quantify without training a custom classification model. Similarly, textual descriptions might reveal attributes that are hard to quantify. Customer reviews, for instance, might highlight that certain headphones are ideal for listening to audiobooks, revealing a benefit not captured by observed attributes. Our approach allows researchers to estimate substitution patterns along these otherwise unobserved dimensions.
A natural consequence of using text and images is that our approach scales more easily across categories than standard attribute-based models. For instance, Döpper et al. (2024) estimate demand estimation across a large set of categories and note that incorporating product attributes would be difficult because it would require them to decide which attributes are relevant in each category and collect data on those attributes. Our approach circumvents the need to collect category-specific attributes, making it easier to estimate demand across many categories. This scalability may also be valuable for online retailers like Amazon or Alibaba looking to optimize prices across hundreds of categories. In Section 4, we further illustrate this point by applying our approach across 40 Amazon categories.
3 Choice Experiment
We begin by applying our method to data from a choice experiment. In this experiment, we randomize product prices and positions, generating clean variation that helps estimate substitution patterns. We also collect data on consumers’ second choices, which are rarely available in observational datasets. These data allow us to assess how well our model recovers substitution patterns by testing its ability to predict second choices counterfactually relative to alternative models.
@@ 3.1 Experiment Design
We recruited participants from the online platform Prolific and asked each participant to complete two choice tasks via a Qualtrics survey.
In the first choice task, we asked each participant to choose one book from a set of ten alternatives (Figure 1). These ten books were chosen from Amazon’s bestseller lists (Figure 2). We chose them from three different genres—Mystery, Fantasy, and Self-Help—anticipating that genre would predict substitution patterns and could thus be incorporated in attribute-based demand models. We instructed each participant to choose the book they would purchase if faced with this selection in a real bookstore.
Figure 1 shows an example of a choice task, displaying the top portion of the page as it appeared to participants. For each book, we showed participants the cover image, title, author, publication year, genre, number of pages, description, and five customer reviews, all collected from Amazon product pages. We presented books in a random order to each participant and drew each book’s price from a discrete uniform distribution ranging from \$3 to \$7.
@@
After a participant selected a book from the list of ten options, we removed this book from the list. The participant then proceeded to a second choice task, where they chose from the remaining nine books. The positions and prices of the books remained the same as in the first choice task, with the only difference being that the participant’s first choice was no longer available.
Based on power calculations from a pilot experiment, we estimated that we needed around 10,000 participants to determine whether our model outperforms alternative models in predicting second choices. We pre-registered this target sample size. [^11] The pre-registration document for this study can be accessed at https://tinyurl.com/ekjp4pfp. In total, we recruited 10,775 participants between June 14 and June 27, 2024. Following the pre-registered sample selection criteria, we excluded around 14% participants who failed comprehension questions, did not complete the survey, or spent less than one minute on the entire study. After applying these criteria, our final sample consists of 9,265 participants.
@@
Because choice tasks in our experiment were hypothetical and not incentivized, it is important to verify that participants made meaningful selections. In Appendix B, we show that participants did not rush through the survey, responded to changes in book rankings and prices, and made choices consistent with their self-reported genre preferences.
3.2 Model Specifications
We compare our approach against two benchmark models: the plain logit model and a standard random coefficients model that leverages observed product attributes but uses no texts or images. All three models fall into the following framework, where the indirect utility of consumer $i$ from purchasing book $j$ is given by
$u_{ij}=\beta_{i}^{\prime}x_{j}+\theta_{i}^{\prime}PC_{j}-\gamma\cdot\text{rank}_{ij}-\alpha_{i}\cdot\text{price}_{ij}+\delta_{j}+\varepsilon_{ij}.$ (1)
Here $x_{j}$ is a vector of observed book attributes, $PC_{j}$ are principal components extracted from text and image embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in consumer $i$’s choice tasks, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d. taste shocks following a Type I Extreme Value distribution. We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books. This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does. When specifying observed attributes $x_{j}$, we include genre dummies for Mystery and Sci-Fi, publication year, and length in pages—all attributes participants observe in the choice tasks.
We estimate the following three models within this framework: [^12] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns related to product features, captured through observed characteristics or principal components from text and image data.
  1. 1.
    Mixed Logit with Principal Components. We include principal components $PC_{j}$, interacting them with random coefficients $\theta_{i}\sim N(0,\Sigma_{\theta})$, where $\Sigma_{\theta}$ is diagonal. Since the principal components do not vary for a given product over time, the mean of the random coefficients is absorbed by the product fixed effects $\delta_{j}$. In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$. We do not include observed attributes, setting $\beta_{i}=0$. This model mimics a scenario where a researcher has access to unstructured data but does not observe attributes other than price.
  2. 2.
    Mixed Logit with Observed Attributes. We include observed attributes $x_{j}$, interacting them with random coefficients $\beta_{i}\sim N(0,\Sigma_{\beta})$, where $\Sigma_{\beta}$ is diagonal, and again, the mean is absorbed by product fixed effects. In some specifications, we also estimate a random coefficient on price, assuming $\alpha_{i}\sim N(\bar{\alpha},\sigma_{\alpha})$. We do not include principal components $PC_{j}$, setting $\theta_{i}=0$. This model reflects the standard practice of estimating substitution patterns through an interaction of observed attributes with consumer heterogeneity.
  3. 3.
    Plain Logit. We include neither attributes $x_{j}$ not principal components $PC_{j}$, estimating only the uniform price coefficient $\alpha_{i}=\alpha$, the rank coefficient $\gamma$ and product fixed effects $\delta_{j}$. This model provides a baseline where product substitution is solely driven by the average utilities of books and not their correlations.
Comparing plain logit to mixed logit with observed attributes tells us whether adding random coefficients on book attributes improves the estimation of substitution patterns. Similarly, comparing mixed logit with attributes to mixed logit with principal components tells us whether incorporating texts and images into demand estimation enables us to recover substitution patterns as well as, or better than, using product attributes. [^13] To ensure a fair comparison between models, for each model we start optimization from 100 different initial points covering a wide range of parameter values. This helps us maximize the probability of finding the global optimum in each model.
3.2.1 Model Selection
Within each of the three model classes listed in Section 3.2, we perform model selection based on $AIC$ [^14] An advantage of using AIC is that it can be interpreted as an estimator of the expected relative Kullback-Leibler (KL) Information based on the maximized log-likelihood function, corrected for asymptotic bias (Akaike, 1998; Burnham and Anderson, 2004). In the mixed logit model with observed attributes, we consider all possible combinations of random coefficients across four book attributes: price, publication year, length in pages, and genre. Genre is represented by two dummy variables, one each for Mystery and one for Self-Help, making Sci-Fi the baseline genre. We estimate all 16 specifications and select the one with the lowest $AIC$.
In the mixed logit model with principal components, we take the first three principal components from each set of embeddings and evaluate all possible combinations of random coefficients on the three components and price, selecting the specification with the lowest $AIC$ [^15] Increasing the number of principal components beyond three has only a marginal impact on our results and does not affect the in-sample fit or counterfactual performance of our selected model. We repeat this selection process for every text and image model and data type (e.g., USE embeddings from reviews). Lastly, among the selected specifications, we choose the one that minimizes $AIC$.
We do not combine principal components extracted from different embeddings (e.g, mixing texts with images) in our main analysis. As we discuss later in Section 3.4.4, we tested mixing data types in our experimental data but found no improvement in model fit.
3.3 Validation Approach
Prior research often evaluated demand models based on how intuitive the estimated cross-price elasticities look, given researchers’ prior knowledge of the market (Berry et al., 1995; Nevo, 2001). We follow a different approach by leveraging second choice data to compare our approach against alternatives. We estimate each model using only first choices and evaluate how well it predicts second choices. [^16] Conceptually, our validation approach reverses the logic of Berry et al. (2004), who show that second choice data are useful for estimating substitution patterns. By contrast, we omit second choice data from estimation and use them only for model validation.
Since the second-choice data are not used in estimation, our validation approach forces alternative models to confront a difficult counterfactual prediction problem. Second choices reveal which books consumers substitute to when their most preferred option becomes unavailable. Therefore, a model can only predict second choices well if it has accurately captured the true substitution patterns. Predicting second choices is also directly relevant in some applications. For instance, antitrust authorities can use them in merger analysis to calculate diversion ratios—a measure reflecting substitutability and thus the intensity of price competition between products (Conlon and Mortimer, 2021). We discuss this further in Section 3.6.
We evaluate each model as follows. First, we obtain MLE estimates by maximizing the total likelihood of participants’ first choices. Given these point estimates $\hat{\theta}$, we then compute the counterfactual second choice probabilities for each participant conditional on their first choice. Averaging these probabilities across participants, we construct the matrix $\mathbb{P}$, where each element $(j,k)$ represents the probability that a participant whose first choice is $j$ selects $k\neq j$ as their second choice. To assess each model, we compute the $RMSE$ based on element-wise differences between the matrix $\hat{\mathbb{P}}$ predicted by the model and $\mathbb{P}$ computed from empirical probabilities in the experimental data. A lower $RMSE$ indicates better model performance in predicting counterfactual second choices.
3.4 Validation Results
We compare the results of three models described in Section 3.2: (1) the plain logit model, (2) the mixed logit model with observed attributes, and (3) the mixed logit model with principal components. We evaluate these models based on the $RMSE$ of their counterfactual second-choice predictions. Figures 3 and 4 summarize the validation results, whereas Appendix Table A1 reports the $AIC$ and $RMSE$ values for all specifications considered within each model class.
3.4.1 Plain Logit
We first report the results from the plain logit model. As shown in Appendix Table A1, this model achieves a counterfactual $RMSE$ of $0.089$. Because this model assumes independence of irrelevant alternatives (IIA), it cannot capture which close substitutes participants switch to when their first choices become unavailable. As a result, the model predicts relatively uniform second-choice probabilities that do not match the second choice data.
3.4.2 Mixed Logit with Principal Components
Next, Figure 3 presents results from the mixed logit model with principal components. The first panel shows $RMSE$ from the plain logit model, indicated by a vertical dashed line, while other panels display results from our proposed approach, which incorporates principal components obtained from different types of unstructured data and machine learning models. Each marker reflects the second-choice $RMSE$ from the best configuration of random coefficients on the principal components and price chosen by our model selection algorithm (see Section 3.2.1).
Performance varies widely across specifications. While some models perform similarly to the plain logit, others deliver significantly better counterfactual predictions. These differences highlight that researchers should always perform model selection to find the specification that best captures substitution patterns from text and images. In practice, researchers can rely on $AIC$ as the main selection criterion. This works well in our experimental data, as $AIC$ selects the model that also achieves the lowest counterfactual $RMSE$ (see Appendix Table A1).
Figure 3: Counterfactual performance of mixed logit models with principal components in terms of predicting second choices.
Figure 4: Counterfactual performance of attribute-based mixed logit models in terms of predicting second choices.
Images
Figure 3 shows that image-based mixed logit specifications perform well. All models in this class outperform the plain logit model, and the lowest-$AIC$ model, InceptionV3, reduces $RMSE$ by 9.8% relative to the plain logit model (from $0.089$ to $0.080$).
To understand why images predict substitution, consider the book covers used in our study (see Figure 2). Within the same genre, book covers often share similar design elements. For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, feature natural objects like twisted vines and golden roses, and have similar fonts and title placements. Two of these books are by the same author and have nearly identical covers. Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds. Given these visual similarities, we expect image embeddings in our models to capture genre information, which correlates with substitution patterns.
However, we emphasize that image embeddings might be informative about substitution beyond simply revealing the book’s genre. For example, second-choice data shows that participants perceive mystery books as closer substitutes to fantasy books than self-help titles (Appendix Figure A3). Our image embeddings capture this pattern because mystery and fantasy book covers tend to be similar. Both use dark or muted tones with dramatic highlights (e.g., gold, silver, teal, or yellow) to create contrast, which is captured by our image embeddings.
@@ 3.4.2 Mixed Logit with Principal Components
The remaining panels of Figure 3 show the performance of text-based models. We find that counterfactual second-choice predictions improve when we either use more advanced text models or incorporate richer text data.
@@
First, counterfactual performance improves substantially when we move from simple bag-of-words models to the more advanced USE and ST models. While bag-of-words models provide little to no improvement over the plain logit, both USE and ST consistently outperform the plain logit across all types of text data. This may be because textual descriptions contain detailed information about book plots but do not always use the exact same words or phrases. Therefore, extracting substitution patterns from text requires a natural language model that can accurately measure semantic similarity.
Second, counterfactual performance improves as we include more text data into the demand model. For instance, the USE model achieves a second-choice $RMSE$ of $0.080$ with book titles, $0.075$ with descriptions, and $0.069$ with reviews. This indicates that customer reviews provide more information about substitution patterns than descriptions, which, in turn, are more informative than titles.
@@
These results make intuitive sense. Although book titles are brief, they often contain subtle genre cues. For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin). Book descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals. Reviews go a step further: consumers summarize plots while also expressing opinions. When reviewing mystery novels, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings.
As with images, texts can provide information about substitution beyond simply revealing genres. For example, book descriptions and reviews make it clear that the self-help books Dopamine Detox and Don’t Believe Everything You Think both discuss how modern distractions and mental habits negatively affect well-being. In contrast, another self-help book, The Art of Letting Go, focuses on letting go of painful memories and emotions. These differences may explain why our text-based models correctly identify the first two books as much closer substitutes, something alternative models fail to detect (see Section 3.5 for details).
[…]
3.4.3 Mixed Logit with Observed Attributes
Figure 4 presents the results for the mixed logit model with observed attributes. We compare it against two benchmarks: the plain logit model and the lowest-$AIC$ text-based model identified earlier. Several specifications significantly outperform the plain logit model. The lowest-$AIC$ specification is a mixed logit model with random coefficients on pages, year, and genre. [^17] To be precise, this model achieves the same $AIC$ as two other mixed logit models (i.e., “Price, Pages & Year” and “Pages & Year”), but all three specifications generate identical second-choice predictions, and thus, achieve the same $RMSE$. By capturing similarities in observed attributes, this model breaks the IIA pattern and generates more accurate second-choice predictions. Consequently, it reduces $RMSE$ by 12.9% compared to plain logit (from $0.089$ to $0.077$).
3.4.4 Summary
Taken together, these findings illustrate the value of our approach, showcasing its ability to extract information about substitution patterns from text and images. We deliberately chose books for this study, thinking that observed attributes like genre would strongly predict substitution. Yet, we find that many text-based models achieve comparable $RMSE$ to that of attribute-based models. Moreover, the best-fitting text-based model reduces $RMSE$ by an additional 10% relative to the best attribute-based model. Thus, incorporating text improves counterfactual predictions beyond what is achievable with observed attributes.
@@ 3.4.4 Summary
One might wonder whether texts, images, and attributes provide distinct information about substitution patterns. If so, the best approach to estimating demand would be to incorporate all three data types. To explore this, we start with our selected review-based specification and examine whether model fit can be improved by adding (1) any combination of observed product attributes or (2) any combination of principal components from ResNet, the lowest-$AIC$ image model. In both cases, we find no fit improvements. This result suggests information across the three data types is highly correlated, with text providing the strongest signal of substitution.
@@
Don’t Believe (S) 0.353 Please Tell Me (M) 0.182 Don’t Believe (S) 0.221 Don’t Believe (S) 0.259
Art of Letting Go (S) 0.249 The Inmate (M) 0.152 The Inmate (M) 0.153 Art of Letting Go (S) 0.176
Please Tell Me (M) 0.112 The Housemaid (M) 0.141 Please Tell Me (M) 0.152 Please Tell Me (M) 0.129
The Inmate (M) 0.094 Don’t Believe (S) 0.137 Art of Letting Go (S) 0.135 The Inmate (M) 0.125
The Housemaid (M) 0.057 Serpent & Wings (F) 0.095 The Housemaid (M) 0.132 The Housemaid (M) 0.100
Serpent & Wings (F) 0.042 Art of Letting Go (S) 0.094 Court of Ravens (F) 0.101 Court of Ravens (F) 0.074
Court of Ravens (F) 0.034 Court of Ravens (F) 0.080 Serpent & Wings (F) 0.057 The Ritual (M) 0.049
The Ritual (M) 0.031 Ashes & Star (F) 0.065 The Ritual (M) 0.028 Serpent & Wings (F) 0.044
Ashes & Star (F) 0.030 The Ritual (M) 0.054 Ashes & Star (F) 0.022 Ashes & Star (F) 0.043

### v3

@@ 2 Proposed Approach
There are several reasons why texts and images may predict substitution. Consumers may choose based on how a product looks and how it is described by sellers or other consumers. Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect attributes that drive substitution. Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly. Similarly, photos may showcase design features like color and casing style. Our approach extracts this information and uses it to capture substitution.
Given this interpretation, embeddings correlate with substitution patterns because they capture choice-relevant attributes, but the relationship need not be causal: changing product images or descriptions will not necessarily alter substitution patterns. Consequently, we cannot study questions like optimal product design or positioning unless we further unpack the link between embeddings and substitution, which is outside the scope of this paper. By contrast, our method is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations. We provide a more detailed discussion of the counterfactuals enabled by our approach in Section 5.2.
Advantages Over Standard Methods
Our approach has several advantages relative to the standard attribute-based methods. First, researchers typically select a limited set of attributes based on their prior knowledge of the market, often those that are easiest to quantify, or rely on attributes supplied by data providers. Our approach avoids these subjective choices by automatically extracting information about product substitution from unstructured data. Second, this approach captures visual design and functional benefits, which may drive substitution but are difficult to capture through observed attributes. Lastly, we circumvent the need to collect category-specific attributes, which makes our approach more scalable. This is valuable for researchers seeking to estimate demand and study competition and pricing across many product categories.
3 Validation with Experimental Data
We apply our method to data from a choice experiment to show it recovers substitution patterns more effectively than standard attribute-based methods. A key challenge is that we usually do not directly observe substitution patterns in the data, and thus do not have a “ground truth” to validate demand models. To address this, we design and implement a choice experiment that measures second-choice diversion ratios, which directly reflect substitution.
@@ 3.1 Experiment Design
Choice Tasks.
Each participant completes two choice tasks. In the first choice task, a participant chooses one book from a set of ten alternatives. We instruct participants to choose books they would purchase if faced with this selection in a real bookstore. We limit the number of options to ten to ensure participants pay attention to most options, which allows us to abstract away from limited consideration. After a participant selects a book from the list of ten options, we remove this book from the list. The participant then proceeds to a second choice task, where they choose among the remaining nine books.
Figure 1 shows the choice task as presented to participants. To give all models a fair chance at capturing substitution, we display attributes (author, year, genre, pages), cover images, and texts (plot description and five reviews), collected from Amazon product pages. We take all books from Amazon’s bestseller lists, sampling them in a way that does not bias our model comparisons in favor of any particular specification (see Appendix B for details). Lastly, we randomize books’ rankings and prices across participants, keeping them fixed across choice tasks for the same participant. [^14] Prices were drawn from a discrete uniform distribution ranging from \$3 to \$7. This double randomization generates clean variation for estimating substitution patterns (Berry and Haile, 2024). [^15] Although all participants technically see the same set of ten books in the first choice task, they rarely choose books that are displayed near the bottom of the list. Therefore, by randomizing the order in which books are displayed to participants, we vary their effective choice sets.
The book category is well-suited for our analysis because participants are likely to consider both structured attributes (e.g., genre) and unstructured information (e.g., plot descriptions). It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for a detailed discussion).
@@
Recruitment and Sample selection.
We recruited 10,775 participants from the online platform Prolific between June 14 and June 27, 2024. This sample size is close to our pre-registered target of 10,000, determined from power calculations in a pilot experiment. [^16] The pre-registration document can be accessed at https://tinyurl.com/ekjp4pfp. At the time of pre-registration, we planned to estimate a different choice model—the pairwise combinatorial logit (PCL) model from Koppelman and Wen (2000)—which generalizes the logit model by allowing utility correlations across product pairs. We switched to a mixed logit model at a later stage because we found it to perform better. We adhered to our pre-registered protocol for all key aspects of survey design, including sample size, sample selection, and choice task construction. Following the pre-registered sample selection criteria, we excluded 14% participants who failed comprehension questions, did not complete the survey, or spent less than one minute on the study, leaving a final sample of 9,265 participants.
Because the choice tasks were hypothetical and not incentivized, it is important to verify that participants made meaningful selections. In Appendix C, we show that participants did not rush through the survey, responded to changes in book rankings and prices, and made choices consistent with their self-reported genre preferences.
3.2 Model Specifications and Estimation Results
We compare our approach against two benchmark models: the plain logit model and a standard logit model with random coefficients model on attributes. All three models fall into a framework where the indirect utility of participant $i$ from book $j$ is
$u_{ij}=\beta_{i}^{\prime}x_{j}+\theta_{i}^{\prime}PC_{j}+\gamma\cdot\text{rank}_{ij}+\alpha_{i}\cdot\text{price}_{ij}+\delta_{j}+\varepsilon_{ij}.$ (1)
Here $x_{j}$ are observed book attributes: genre dummies, publication year, and length in pages—all attributes participants observe in the choice tasks; $PC_{j}$ are principal components extracted from embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in participant $i$’s choice task, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d. taste shocks following a Type I Extreme Value distribution. We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books. [^17] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns using observed characteristics or principal components. This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does.
We estimate three demand models:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    Plain Logit. We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$ and fixed effects $\delta_{j}$.
Table 1 compares in-sample $AIC$. The best mixed logit with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit. The best model with unstructured data, Reviews USE, puts random coefficients on the first two principal components. This model fits better than any other considered specification, reducing $AIC$ by $24.8$ relative to plain logit and by $8.8$ relative to the best attribute-based logit. While this superior fit is reassuring, it does not guarantee better counterfactual performance. We therefore turn to predicting second choices counterfactually.
[…]
Plain Logit None 41006.7 0.0
Mixed Logit with Attributes Pages and Year 40990.7 -16.0
Mixed Logit with Images (InceptionV3) Price, PC1, and PC6 40990.5 -16.2
Mixed Logit with Titles (ST) PC1 and PC5 40992.3 -14.4
Mixed Logit with Descriptions (USE) PC1 and PC5 40986.4 -20.3
Mixed Logit with Reviews (USE) PC1 and PC2 40981.9 -24.8
Table 1: Comparison of models in terms of in-sample AICAIC on first choices. The second column shows variables that have random coefficients in the selected specification. The last column shows the $AIC$ reduction relative to plain logit.
3.3 Validation Using Second-Choice Data
To evaluate demand models, we let them confront a difficult counterfactual prediction problem—predicting second choices, i.e. which books participants switch to when their most preferred option becomes unavailable. A model can only predict second choices well if it has accurately captured the true substitution patterns. Predicting second choices is also directly relevant in antitrust where second-choice diversion ratios are often used to measure substitutability and the intensity of price competition (Conlon and Mortimer, 2021).
Operationally, we first estimate each model via Maximum Likelihood Estimation (MLE) using first choices only. We then use the estimated model to predict counterfactual second-choice diversion ratios, $\hat{s}_{j\rightarrow k}$, defined as the probability that the participant chooses product $k$ in the second choice task conditional on having chosen book $j$ in the first choice task. We compare these predictions with diversions observed in the data, $s_{j\rightarrow k}$, computing $RMSE$ as:
$RMSE=\sqrt{\frac{1}{J(J-1)}\sumop\displaylimits_{j}\sumop\displaylimits_{k\neq j}\left(s_{j\rightarrow k}-\hat{s}_{j\rightarrow k}\right)^{2}}$ (2)
Since we do not use second-choice data in estimation, lower $RMSE$ indicates that the model performs better at counterfactual predictions of second choices. Appendix D provides further details on how we compute diversions $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$.
Figure 2: Comparison of models in terms of counterfactual RMSERMSE on second choices. The two benchmarks in the first panel are the plain logit without random coefficients and the best-fitting mixed logit with random coefficients on observed attributes. The remaining specifications correspond to mixed logit models with random coefficients on principal components extracted from image or text embeddings.
Figure 2 summarizes the validation results, while Appendix Table A1 reports $AIC$ and $RMSE$ values for all estimated specifications. The top panel in Figure 2 shows two benchmarks: the plain logit and the lowest-$AIC$ mixed logit with attributes. The mixed logit with attributes reduces RMSE by 11.7%, while our best-fitting Review USE model reduces RMSE by 23% relative to plain logit, significantly outperforming the mixed logit with attributes. This result illustrates the value of our approach. We chose books for this study expecting observed attributes like genre to predict substitution well. Yet, by using unstructured data, we can match the counterfactual performance of the mixed logit with attributes and even further reduce $RMSE$ by 14%. This shows that our approach recovers substitution patterns better than standard attribute-based methods in this dataset.
Although the model with product reviews performs best, in general, performance varies widely across specifications. Below, we discuss why some specifications recover substitution patterns better than others.
Images
All four image models outperform the plain logit model, with the lowest-$AIC$ model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit. To understand why images predict substitution, consider the book covers used in our study (see Figure 3). Within the same genre, book covers often share similar design elements. For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, and feature natural objects like twisted vines and golden roses. Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds. Thus, image embeddings partly encode books’ genres, which correlate with substitution patterns.
@@ 3.3 Validation Using Second-Choice Data
All text models outperform the plain logit, and their relative performance highlights several notable patterns.
@@
First, performance tends to improve when we move from simple bag-of-words models to the more advanced USE and ST models. This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always describe them in the same words or phrases. Therefore, extracting substitution patterns from text requires natural language models that can accurately measure semantic similarity.
@@
Second, performance improves as we include richer text data. For instance, the USE model reduces $RMSE$ by 4.3% with book titles, 17% with descriptions, and 23% with reviews relative to the plain logit. This makes intuitive sense. Although book titles are brief, they often contain subtle genre cues. For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin). Descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals. Reviews go even further: consumers summarize plots while also expressing opinions. When reviewing mystery novels, for instance, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings. Thus, it is unsurprising that reviews provide more information about substitution than descriptions, which, in turn, are more informative than titles.
Extension: Combining Data Types
@@
If text, images, and attributes provides distinct signals of substitution, combining them might improve fit and counterfactual performance. To explore this, we start from the selected Review USE model and attempt to extend it in two ways: (1) by adding a combination of observed product attributes with random coefficients, or (2) by adding a combination of principal components from InceptionV3, the lowest-$AIC$ image model. Neither extension improves $AIC$ or $RMSE$. This result suggests information across data types is highly correlated, with text providing the strongest signal of substitution.
What do Principal Components Capture?
To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-fitting Review USE model. These principal components align with intuitive substitution patterns: the first one (horizontal axis) separates non-fiction on the left from fiction on the right, while the second (vertical axis) further distinguishes science fiction from mystery. Crucially, the variation in these principal components goes beyond separating genres. For example, The Housemaid and The Inmate are by the same author, and The Serpent & The Wings and The Ashes & The Star Cursed King are from the same book series. The two principal components detect this similarity from consumers’ reviews of these books. As evident from $RMSE$ comparisons in Figure 2, this additional variation helps recover substitution patterns better and improves counterfactual predictions.
Figure 4: Book locations in the space of selected principal components (Review USE model).
@@
Don’t Believe (S) 0.353 Please Tell Me (M) 0.171 Don’t Believe (S) 0.220 Don’t Believe (S) 0.266
Art of Letting Go (S) 0.249 The Inmate (M) 0.147 Please Tell Me (M) 0.153 Art of Letting Go (S) 0.169
Please Tell Me (M) 0.112 Don’t Believe (S) 0.143 Art of Letting Go (S) 0.142 Please Tell Me (M) 0.134
The Inmate (M) 0.094 The Housemaid (M) 0.139 The Housemaid (M) 0.134 The Inmate (M) 0.123
The Housemaid (M) 0.057 Art of Letting Go (S) 0.106 The Inmate (M) 0.131 The Housemaid (M) 0.101
Serpent & Wings (F) 0.042 Court of Ravens (F) 0.096 Court of Ravens (F) 0.101 Court of Ravens (F) 0.070
Court of Ravens (F) 0.034 Serpent & Wings (F) 0.088 Serpent & Wings (F) 0.060 Serpent & Wings (F) 0.057
The Ritual (M) 0.031 The Ritual (M) 0.057 The Ritual (M) 0.031 The Ritual (M) 0.043
Ashes & Star (F) 0.030 Ashes & Star (F) 0.054 Ashes & Star (F) 0.027 Ashes & Star (F) 0.037